Two vectors p and q are given by:

p = 2i + j - K

q = i - 3j + 2k

Find:

a) p + q

b) Modulus (p + q)

c) Modulus (p) + Modulus (q)

d) (p) x (q)
Answer #1

a) p + q = (2i + j - k) + (i - 3j + 2k)

= 2i + 1j - 1k + 1i - 3j +2k

= i(2+1) + j(1-3) + k(-1+2)

= i(3) + j(-2) + k(1)

= 3i - 2j + 1k

= 3i -2j + k (Ans)

b) Modulus (p+q) = [ (3)2 + (-2)2 + (1)2 ]1/2

= [ 9 + 4 + 1]1/2

= [14]1/2

= 3.7416 (Ans)

c) Modulus (p) + Modulus (q) = [ (2)2 + (1)2 + (-1)2]1/2 + [ (1)2 + (-3)2 + (2)2]1/2

= [ 4 + 1 + 1]1/2 + [ 1 + 9 + 4]1/2

= [6]1/2 + [14]1/2

= 2.4495 + 3.7416

= 6.1911 (Ans)

d) Before we move into multiplication, some things to remember,

i * i = 0, i * j = 1k, i * k = -1j, j*i = -1k, j * j = 0, j * k = 1i, k*i = 1j, k * j = -1i, k*k = 0

(p) * (q) = (2i + j -k) * (i - 3j + 2k)

= 2i * (i - 3j + 2k) + j * (i - 3j +2k) - k * (i - 3j +2k)

= (2 * i * i )+  (2 * -3 * i * j) + (2 * 2 * i * k) + ( j * i) + ( 1 * - 3 * j * j) + ( 1 * 2 * j * k) + ( -1 * 1 * k * i) + (-1 * -3 * k * j) + (-1*2 * k * k)

= (2 * 0) + ( -6 * 1k) + ( 4 * -1j ) + (-1k) + ( -3 * 0 ) + ( 2 * 1i) + ( -1 * 1j) + ( 3 * -1i) + ( -2 * 0)

= (-6k) + ( -4j) + (-1k) + (2i) + (-1j) + (-3i)

= -6k - 4j -k + 2i -j -3i

= -6k -k - 4j -j +2i - 3i

= -7k - 5j -1i

= -1i -5j - 7k (Ans)